Dx/y^2+z^2=dy/-xy=dz/-xz


1
Expert's answer
2021-05-07T14:19:12-0400

We have to solve the given differential equation,


"\dfrac{dx}{y^2+z^2} = \dfrac{dy}{-xy} = \dfrac{dz}{-xz}"


Firstly, we have to solve last two equations:

then,


"\dfrac{dy}{-xy} = \dfrac{dz}{-xz}"


"\dfrac{dy}{y} = \dfrac{dz}{z}"


"logy = logz+logC"


"y = zC"


Now, we will solve first and last equations:


"\dfrac{dx}{y^2+z^2} = \dfrac{dz}{-xz}"


Putting "y = Cz"


"\dfrac{dx}{z^2C^2+z^2} = -\dfrac{dz}{xz}"


"\dfrac{dx}{z(C^2+1)} = -\dfrac{dz}{x}"


"xdx = -(C^2+1)dz"


Integrating both sides


"\int xdx = - \int (C^2+1)dz"


"\dfrac{x^2}{2} = -(C^2+1)z + C_1"